I need to know if this is identity or contradiction.
x ( x - 3 ) = x^2 - 2x + 1 - ( 5 + x )

Answer by stanbon(75887)

x ( x - 3 ) = x^2 - 2x + 1 - ( 5 + x )
-----
x^2-3x = x^2-2x+1-5-x
-----
x^2-3x = x^2 -3x -4
----------
0 = -4
--------
It is a contradiction.
===================
Cheers,
Stan H.
===================